A 740N student is standing on a frictionless frozen pond with a radius of 5.2 meters. He throws a 3.2 kg textbook toward the shore at 5.6m/s. Gravity = 9.81m/s^2. How long does it take the student to reach the shore in seconds?
Answer #1

Mass of student

M=740/9.81 =75.433 Kg

By Conservation of momentum

MV+mv =0

75.433*V+3.2*5.6=0

V=-17.92/75.433=-0.23756 m/s

V=0.23756 m/s (in opposite direction to book)

Time taken by student to reach the shore is

t=d/V =5.2/0.23756

t=21.89 s =21.9 s